OpenStudy (anonymous):

Determine whether y varies directly with x. If so, find the constant of variation k and write the equation. x | y 7 | 11 8 | 13 9 | 15 10| 17

OpenStudy (anonymous):

not varing directly as if we find k=7/11 k=8/13 k=9/15 k is not constant so x and y do not vary directly

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

welcome